# An Algorithmic Approach to Finding Degree-Doubling Nodes in Oriented Graphs

### v7

@@ A Graph Level Order (GLOVER) on a set $S$ can be defined as follows:
[…]
@@
Interior Degree at least 1 Suppose we have an oriented graph $G$, represented by a GLOVER. Suppose further that $v_{0}$ is a minimum out-degree node for $G$ satisfying the decreasing sequence property. In an attempt to prevent $v_{0}$’s degree from doubling in $G$, every node $x$ in the neighboring container $C_{1}$ must have an interior degree of at least 1. That is, for $x\in C_{1}$,
@@
By definition, a node satisfies the decreasing sequence property if $|N^{+}(x)|>N^{++}(x)|$ If $x$ has $d^{+}(v_{0})$ exterior arcs, this would make $v_{0}$’s degree double because $x$’s neighbors are $v_{0}$’s second neighbors. To avoid this doubling, $x$ must connect to at least one interior arc. This means that the interior degree of $x$ with respect to $v_{0}$ must satisfy
@@
This is true for any neighbor $x\in C_{1}$. This ensures that every node in $C_{1}$ has an interior degree of at least 1. Thus, for all $x\in C_{1}$, $|int(v_{0},x)|\geq 1$. ∎
@@
Exterior Less than Degree Suppose we have an oriented graph $G$, represented by a GLOVER. In an attempt to prevent a degree-doubling node, if $x_{i}$ in container $C_{i}$ has the decreasing sequence property, then for all $y\in N^{+}(x_{i})|\text{ext}(x,y)|<d^{+}(x)$.
@@
The results presented in 5.4 (Interior Degree At Least 1) and 5.5 (Exterior Less than Degree) are closely related to Daamouch’s [7] Lemma 1.1. There, he demonstrated that a minimum-degree vertex with no cycles in its induced graph has a degree that effectively doubles.
Despite the overlap, this paper will be moving in a different direction with the containers and the decreasing sequence property towards an global search solution with the algorithm we are developing.
@@
Weak Decreasing Exteriors Lemma Suppose that $G$ is an oriented graph represented by a GLOVER. Suppose further that $x$ is a node in this graph, with $y\in N^{+}(x)$ such that both $x$ and $y$ satisfy the decreasing sequence property (DSP), and suppose that we are trying to prevent a degree doubling node.
@@
    For all $z\in ext(x,y)$, we have $ext(y,z)\cap N^{+}(x)=\emptyset$, i.e., no back arcs from the second neighbors of $x$ to the first neighbors of $x$ exist.
@@
We begin by recalling the conditions for a node to have the decreasing sequence property (DSP) because it is very important for this proof. A node has the DSP if the size of its out-neighborhood is strictly greater than the size of its second out-neighborhood, i.e., $|N^{+}(x)|>|N^{++}(x)|$. Now, assume the conditions of the lemma hold:
@@
Suppose for contradiction that there exists $z\in\text{ext}(x,y)$ such that $|\text{ext}(y,z)|\geq|\text{ext}(x,y)|$. The node $y$ has a degree that can be broken down into two components, its interior degrees and its exterior degrees. By the assumption that $y$’s interior degree doubles, if $y$’s exterior degree were to double, it would lead to a contradiction. If there exists a node $z\in\text{ext}(x,y)$ such that $|\text{ext}(y,z)|\geq|\text{ext}(x,y)|$, then $y$’s out-degree will be impacted by $\text{ext}(x,y)$ and $\text{ext}(y,z)$.
@@
Notice the wording of interior degree doubling. That is a separate version of the SSNC, but localized to the containers. The conjecture in this case is can we find a node in each container whose degree doubles. If so, then it gives light to this divide-and-conquer approach, where we partition the containers, solve them individually, and then combine them again for this lemma 5.6 (Weak Decreasing Exteriors Lemma).
The more popular representations of graphs are adjacency matrices, arraylists, and edge lists. These suffer from the same limitations of graphs as mentioned above; they scatter the nodes along the two-dimensional plane without attempts to group them into an ordering. The Graph Level Order (GLOVER) improves this by partitioning the nodes in a reasonable way. For the SSNC, that reasonable way was the distance metric. To represent graphs in a GLOVER, we can use any data serialization that can encode graphs, including JSON, YAML, and XML. There needs to be three types of data: containers, nodes, and arcs. The containers will be given IDs. The nodes will then be assigned to their unique containers. Arcs will also be determined to be interior, exterior, or back, depending on the endpoints of the nodes of the arcs.
@@
With this container representation of the SSNC, by way of the GLOVER data structure, we can begin to visualize elements of these lemmas. For example, the lemma 5.6 (Weak Decreasing Exteriors Lemma) can be visualized by example 5.4. This shows a series of ordered containers. Each container has fewer nodes. The fewer nodes implies that the container circle can be drawn with a smaller radius, giving light to the decrease in size.
@@
Now, lemma 5.4 (Interior Degree at least 1) says that each person in container 1 must follow another person in container 1. What this is saying is that, inside this first container, we will have one cycle of users where the messages are traversed about or a number of disjoint cycles where messages are spread among themselves. Considering example 5.5, which takes the former of these approaches. These cycles in the first container are what stops the user Andi from being an influencer.
This brings us to the next container. Lemma 5.6 (Weak Decreasing Exteriors Lemma) lets us know that this container will have 5 people—Hugo, Isla, Jade, Kai, and Lei. If any person in this container followed less than two other people in this same container, then they would need to follow at least five people in the next container (container 3). From this we would conclude that a person in container 1, whose indirect sphere of influence is this person’s sphere of influence, would have its first degree double and its second degree double, implying that its overall degree doubles.
@@
$\Rightarrow$ BellChadDawnElseFayeGusHugoIslaJade $\Rightarrow$ KaiLeiMiaNovaOwenPatQinRyan $\Rightarrow$ SageTateUgoVeraWadeXiaYoel $\Rightarrow$ H
@@ 6 Back Arcs
If we look at Example 4.2 again, we see that $N^{+}(v_{0})=\{v_{1},v_{2},v_{3}\}$, $|int(v_{0},v_{1})|=|\{v_{2}\}|=1$ since $(v_{1},v_{2})\in G$. Similarly, we see that $|ext(v_{0},v_{1})|=|\{v_{4},v_{5}\}|=2$ since $(v_{1},v_{4})\in G$ and $(v_{1},v_{5})\in G$ and $(v_{0},v_{4})\in G$ and $(v_{0},v_{5})\in G$.
@@
Figure 9: Graph representation of back arcs (shown in red) and forward-directed arcs (in blue) between containers, illustrating how these arcs impact connectivity patterns in an oriented graph.
@@
(Regular Interior Degrees) Let $G$ be an oriented graph represented by a GLOVER. Let $v_{0}$ be a minimum degree node of $G$. Suppose that the containers of $G$ are $C_{i}$. In an attempt to prevent a degree-doubling node, suppose that two nodes $u$ and $v$ in a container $C_{i}$ have a common parent $v_{p}$ in the preceding container $C_{i-1}$, and that $v_{p}$’s interior degree doubles. If the interior degrees of $u$ and $v$ differ by at least 1, then the exterior degree of $v_{p}$ in container $C_{i-1}$ will double as well.
@@
As a result, the exterior degree of $v_{p}$ will double. Because $v_{p}$ has already had its interior degree doubled, it will have its overall degree double
Thus, the different interior degrees of $u$ and $v$ in container $C_{i}$ leads to a contradiction, so the claim holds. ∎
@@
Table 1: Consequences of Back Arcs in 5.5
@@
No Back Arc to Early Container Let $G$ be an oriented graph represented by a GLOVER in an attempt to prevent a degree-doubling node. If a node $u_{i}$ in a container $C_{i}$ has its first (earliest) back arc to a node $v_{k}$ in container $C_{k}$ where $k<i-1$, then the size of the exterior neighbors $\text{ext}(u_{i},v_{k})$ is at least the degree of $v_{k}$. Furthermore, if the interior degree of $u_{i}$ doubles, this will cause the total degree of $u_{i}$ to double as well.
@@
Back Arcs Special Case Let $G$ be an oriented graph represented by a GLOVER, with containers $C_{1},C_{2},\ldots,C_{n}$ defined based on the distance from the minimum degree node $v_{0}$ in order to prevent a degree doubling node.
@@
First we must notice that the order within the container system, which defines exterior arcs as going to the next container (i.e., from container $C_{i}$ to container $C_{i+1}$) and not two containers down (i.e., not from container $C_{i}$ to container $C_{i+2}$), explicitly forbids $x$ having an exterior neighbor $z$ in the container $C_{i+2}$ where $x$ is in container $C_{i}$. This gives the only possibility for $z$ as being an interior neighbor of $x$.
Consequently, the arc $w,z$ is an early back arc as we saw in 6.2 (No Back Arc to Early Container). These arcs can lead to double counting. An out-neighbor of $w$ might be an out-neighbor of $x$ and a second out-neighbor of $y$. This implies that under the assumption that $w$’s interior degree doubles, $w$’s overall degree will double through that early back arc.
@@
First Back Arc Between Consecutive Containers in an Oriented Graph: Let $G$ be an oriented graph with minimum degree node $v_{0}$, with minimum degree $d$. Suppose that we represent this graph by a GLOVER. Suppose we are attempting to prevent a degree-doubling node and that all arcs point forward (i.e., from container $C_{j}$ to container $C_{j+1}$ for $j<i$) with no back arcs between containers $C_{1},C_{2},\dots,C_{i-1}$. If the first back arc occurs from container $C_{i}$ to $C_{i-1}$, then it will introduce an arc violating $G$’s oriented property.
@@
Lemma 6.4 (No Back Arc to Early Container) is a special case where we can apply Lemma 6.1 (Regular Interior Degrees). From this, we can see that a back arc to a parent container violates the oriented nature of the graph.
@@
No Back Arcs Non-Existence of Back Arc In an oriented graph $G$, represented by a GLOVER in an attempt to prevent a degree-doubling node, with a minimum degree node $v_{0}$ and containers $C_{1}$, $C_{2}$, $\dots$, and $C_{k}$, there cannot exist back arcs between a node $u_{i}\in C_{i}$ and a node $v_{j}\in C_{j}$, where $j<i$.
@@
This was proved in the lemmas 6.2 (No Back Arc to Early Container) and 6.4 (First Back Arc Between Consecutive Containers). ∎
[…]
Figure 10: Illustration of the proof of 7.2 (Interior Degrees Double). This is another illustration of the Interior Degree Doubling. Each node inside a container represents a node whose degree is doubled by the inductive hypothesis. This process cannot continue forever, so we must reach the last node, and the parent of this last node in the sequence is the node whose exterior degree will double.
[…]
Figure 11: Illustration of the proof of 7.2 (Interior Degrees Double). Each circle represents a new container; by the I.H., there exists a $y$ whose interior degree doubles, leading to decreasing exteriors.

### v8

@@ 6 Using the Data Structure
0
$\Rightarrow$ 1111222233334444d $\Rightarrow$ d + $1$d + $2$d + 33332*d-1 $\Rightarrow$
Definition 6.1.
For a node $x\in V(G)$ in an oriented graph $G$, we say that $x$ has the decreasing sequence property if the size of its first neighbors is strictly larger than the size of its second neighbors, i.e., $|N^{+}(x)|>|N^{++}(x)|$.
The SSNC asks for a node whose second neighborhood is at least as large as its first. Our strategy is to prove the conjecture by contradiction. We assume the decreasing sequence property holds for all nodes in $G$. The definition of the decreasing sequence property is the negation of what the conjecture asks for. This assumption allows us to shift our focus from directly searching for a degree-doubling node to analyzing the graph’s global composition. One notable change is that the decreasing sequence property calls for the second neighborhood to be strictly less than the first. This could set up a decreasing exterior condition, which could prove strong. We anticipate that this exploration will reveal constraints on the graph’s arrangement that will ultimately lead to a contradiction.
Two competing new ideas have been presented: the decreasing sequence property (DSP) and the minimal degree node. A crucial starting point for our approach and degree values is a minimal degree node. It acts as the anchor by being chosen as the beginning point. The remaining nodes are currently being forced into the attributes that will surface in the counterexample construction by the minimal degree node. It acts as a catalyst, influencing other nodes and setting up the conditions necessary for the subsequent analysis.
In contrast, the DSP introduces constraints on connectivity. This concept is akin to a bottleneck. A bottleneck in graph theory represents a point that limits flow or connectivity within the network[6], a role precisely filled by the DSP. Thus, the minimum degree node and the DSP together function as a catalyst and bottleneck, respectively.
@@
In this work, we assume that the input graph is oriented and unweighted. We will assume that all nodes in this graph have the decreasing sequence property. What that means is that every node’s second neighborhood is strictly smaller than its first neighborhood. In the case of Seymour’s Second Neighborhood Conjecture (SSNC), this is still the setup for a proof by contradiction. In the previous section, we showed that any unweighted oriented graph without back arcs could be represented in a GLOVER data structure. Instead of allowing nodes to be placed anywhere across the plane, we will take advantage of this structure here. The nodes will be partitioned into containers with interior arcs to the current container and exterior arcs to the next container. For the remainder of this paper, when we speak of oriented graphs, the underlying assumption is that it is an oriented graph in a GLOVER data structure.
Lemma 6.2.
Interior Degree at least 1 Suppose we have an oriented graph $G$. Suppose further that $v_{0}$ is a minimum out-degree node for $G$ satisfying the decreasing sequence property. In an attempt to prevent $v_{0}$’s degree from doubling in $G$, every node $x$ in the neighboring container $C_{1}$ must have an interior degree of at least 1. That is, for $x\in C_{1}$,
@@
By definition, a node satisfies the decreasing sequence property if
$|N^{+}(x)|>N^{++}(x)|$
If $x$ has $d^{+}(v_{0})$ exterior arcs, this would make $v_{0}$’s degree double because $x$’s neighbors are $v_{0}$’s second neighbors.
To avoid this doubling, $x$ must connect to at least one interior arc. This means that the interior degree of $x$ with respect to $v_{0}$ must satisfy
@@
This is true for any neighbor $x\in C_{1}$. This ensures that every node in $C_{1}$ has an interior degree of at least 1.
Thus, for all $x\in C_{1}$, $|int(v_{0},x)|\geq 1$. ∎
@@
Exterior Less than Degree
Suppose we have an oriented graph $G$. If $x_{i}$ in container $C_{i}$ has the decreasing sequence property, then for all $y\in N^{+}(x_{i})|\text{ext}(x,y)|<d^{+}(x)$.
@@
The results presented in 6.2 (Interior Degree At Least 1) and 6.3 (Exterior Less than Degree) are closely related to Daamouch’s [7] Lemma 1.1. There, he demonstrated that a minimum-degree vertex with no cycles in its induced graph has a degree that effectively doubles.
Despite the overlap, this paper will be moving in a different direction with the containers and the decreasing sequence property towards a global search solution with the algorithm we are developing.
@@
Weak Decreasing Exteriors Lemma Suppose that $G$ is an oriented graph. Suppose further that $x$ is a node in this graph, with $y\in N^{+}(x)$ such that both $x$ and $y$ satisfy the decreasing sequence property (DSP), and suppose that we are trying to prevent a degree doubling node.
@@
    For all $z\in\text{ext}(x,y)$, we have $\text{ext}(y,z)\cap N^{+}(x)=\emptyset$, i.e., no back arcs from the second neighbors of $x$ to the first neighbors of $x$ exist.
@@
We begin by recalling the conditions for a node to have the decreasing sequence property (DSP) because it is very important for this proof. A node has the DSP if the size of its out-neighborhood is strictly greater than the size of its second out-neighborhood, i.e., $|N^{+}(x)|>|N^{++}(x)|$.
Now, assume the conditions of the lemma hold:
@@
Suppose for contradiction that there exists $z\in\text{ext}(x,y)$ such that $|\text{ext}(y,z)|\geq|\text{ext}(x,y)|$.
The node $y$ has a degree that can be broken down into two components, its interior degrees and its exterior degrees.
By the assumption that $y$’s interior degree doubles, if $y$’s exterior degree were to double, it would lead to a contradiction.
If there exists a node $z\in\text{ext}(x,y)$ such that $|\text{ext}(y,z)|\geq|\text{ext}(x,y)|$, then $y$’s out-degree will be impacted by $\text{ext}(x,y)$ and $\text{ext}(y,z)$.
@@
Notice the wording of interior degree doubling. That is a separate version of the SSNC, but localized to the containers. The conjecture in this case is, can we find a node in each container whose degree doubles? If so, then it gives light to this divide-and-conquer approach, where we partition the containers, solve them individually, and then combine them again for this lemma 6.4 (Weak Decreasing Exteriors Lemma).
The more popular representations of graphs are adjacency matrices, array-lists, and edge lists. These suffer from the same limitations of graphs as mentioned above; they scatter the nodes along the two-dimensional plane without attempts to group them into an ordering. The Graph Level Order (GLOVER) improves this by partitioning the nodes in a reasonable way. For the SSNC, that reasonable way was the distance metric. To represent graphs in a GLOVER, we can use any data serialization that can encode graphs, including JSON, YAML, and XML. There need to be three types of data: containers, nodes, and arcs. The containers will be given IDs. The nodes will then be assigned to their unique containers. Arcs will also be determined to be interior, exterior, or back, depending on the endpoints of the nodes of the arcs.
@@
With this container representation of the SSNC, by way of the GLOVER data structure, we can begin to visualize elements of these lemmas. For example, the lemma 6.4 (Weak Decreasing Exteriors Lemma) can be visualized by example 6.2. This shows a series of ordered containers. Each container has fewer nodes. The fewer nodes imply that the container circle can be drawn with a smaller radius, giving light to the decrease in size.
Now, lemma 6.2 (Interior Degree at least 1) says that each person in container 1 must follow another person in container 1. What this is saying is that, inside this first container, we will have one cycle of users where the messages are traversed about or a number of disjoint cycles where messages are spread among themselves. Considering example 6.3, which takes the former of these approaches. These cycles in the first container are what stop the user, Andi, from being an influencer.
This brings us to the next container. Lemma 6.4 (Weak Decreasing Exteriors Lemma) lets us know that this container will have 8 people—Kai, Lei, Mia, Nova, Owen, Pat, Qin, and Ryan. If any person in this container followed less than two other people in this same container, then they would need to follow more people in the next container (container 3). From this we would conclude that a person in container 1, whose indirect sphere of influence is this person’s sphere of influence, would have its first degree double and its second degree double, implying that its overall degree doubles.
@@
$\Rightarrow$ BellChadDawnElseFayeGusHugoIslaJade $\Rightarrow$ KaiLeiMiaNovaOwenPatQinRyan $\Rightarrow$ SageTateUgoVeraWadeXiaYoel $\Rightarrow$
@@ 7 Back Arcs
If we look at Example 4.2 again, we see that $N^{+}(v_{0})=\{v_{1},v_{2},v_{3}\}$, $|int(v_{0},v_{1})|=|\{v_{2}\}|=1$ since $(v_{1},v_{2})\in G$. Similarly, we see that $|\text{ext}(v_{0},v_{1})|=|\{v_{4},v_{5}\}|=2$ since $(v_{1},v_{4})\in G$ and $(v_{1},v_{5})\in G$ and $(v_{0},v_{4})\in G$ and $(v_{0},v_{5})\in G$.
@@
(Regular Interior Degrees) Let $G$ be an oriented graph. Let $v_{0}$ be a minimum degree node of $G$. Suppose that the containers of $G$ are $C_{i}$. In an attempt to prevent a degree-doubling node, suppose that two nodes $u$ and $v$ in a container $C_{i}$ have a common parent $v_{p}$ in the preceding container $C_{i-1}$, and that $v_{p}$’s interior degree doubles. If the interior degrees of $u$ and $v$ differ by at least 1, then the exterior degree of $v_{p}$ in container $C_{i-1}$ will double as well.
@@
As a result, the exterior degree of $v_{p}$ will double. Because $v_{p}$ has already had its interior degree doubled, it will have its overall degree doubled.
Thus, the different interior degrees of $u$ and $v$ in container $C_{i}$ lead to a contradiction, so the claim holds. ∎
@@
The assumptions for this letter are that we have one parent node whose interior degree doubles, and it maps the two child nodes that have different interior degrees. What happens is that because these child nodes have different interior degrees, it causes this pair node to have an exterior degree that doubles by the week exterior decreasing lemma.
We spoke previously about how the divide and conquer approach is allowed by the GLOVER data structure. This is now the second lemma that we’ve seen call for interior degrees doubling. That is going to be a staple of this algorithm as we see the divide and conqueror approach called for more often as this interior degree doubling is part of the nodes being called back into the global solution. Remember that interior degrees doubling are local instances of solving the Seymour second neighborhood conjecture on the container itself. Instead of seeking a degree-doubling node, these containers seek an interior degree-doubling node. What remains to be seen is if interior degree doubling is an effective solution. That will show in the next section.
@@
Table 1: Consequences of Back Arcs in 6.3
@@
No Back Arc to Early Container Let $G$ be an oriented graph. If a node $u_{i}$ in a container $C_{i}$ has its first (earliest) back arc to a node $v_{k}$ in container $C_{k}$ where $k<i-1$, then the size of the exterior neighbors $\text{ext}(u_{i},v_{k})$ is at least the degree of $v_{k}$. Furthermore, if the interior degree of $u_{i}$ doubles, this will cause the total degree of $u_{i}$ to double as well.
@@
The proof above demonstrates that a node cannot have a back arc to an early container. To understand this proof, we must first investigate the assumptions. Here, we are assuming two things. That the node $u_{i}$ has its interior degree doubled and that the first back arc from $u_{i}$ is to the node $v_{k}$. As we saw in the proof, $v_{k}$ has more exterior neighbors than you survive, which is what causes $u_{i}$’s degree to double. The question is whether we need the condition of $u_{i}$ interior degree doubling. The weak decreasing exterior lemma tells us that we have a relationship between the exterior degrees of use by and $v_{k}$, not the overall degrees of these nodes. By assigning the interior degree to double, we can leverage this lemma.
Secondly, we look at the notion of the first back arc. The containers are totally ordered. By mapping $u_{i}$ to its first back arc and showing that this back arc cannot exist, we have shown that $u_{i}$ cannot have a first back arc. Because $u_{i}$ cannot have a first back arc, this back arc of this type cannot exist.
@@
Back Arcs Special Case
Let $G$ be an oriented graph, with containers $C_{1},C_{2},\ldots,C_{n}$ defined based on the distance from the minimum degree node $v_{0}$.
@@
First, we must notice that the order within the container system, which defines exterior arcs as going to the next container (i.e., from container $C_{i}$ to container $C_{i+1}$) and not two containers down (i.e., not from container $C_{i}$ to container $C_{i+2}$), explicitly forbids $x$ having an exterior neighbor $z$ in the container $C_{i+2}$ where $x$ is in container $C_{i}$. This gives the only possibility for $z$ as being an interior neighbor of $x$.
Consequently, the arc $w,z$ is an early back arc as we saw in 7.2 (No Back Arc to Early Container). These arcs can lead to double counting. An out-neighbor of $w$ might be an out-neighbor of $x$ and a second out-neighbor of $y$. This implies that under the assumption that $w$’s interior degree doubles, $w$’s overall degree will double through that early back arc.
@@
First Back Arc Between Consecutive Containers in an Oriented Graph: Let $G$ be an oriented graph with minimum degree node $v_{0}$, with minimum degree $d$. Suppose that all arcs point forward (i.e., from container $C_{j}$ to container $C_{j+1}$ for $j<i$) with no back arcs between containers $C_{1},C_{2},\dots,C_{i-1}$. If the first back arc occurs from container $C_{i}$ to $C_{i-1}$, then it will introduce an arc violating $G$’s oriented property.
@@
Lemma 7.4 (No Back Arc Between Consecutive Containers) is a special case where we can apply Lemma 7.1 (Regular Interior Degrees). From this, we can see that a back arc to a parent container violates the oriented nature of the graph.
@@
No Back Arcs Non-Existence of Back Arc In an oriented graph $G$, with a minimum degree node $v_{0}$. Let containers $C_{1}$, $C_{2}$, $\dots$, and $C_{k}$. Then there cannot exist back arcs between a node $u_{i}\in C_{i}$ and a node $v_{j}\in C_{j}$, where $j<i$.
@@
This was proved in the lemmas 7.2 (No Back Arc to Early Container) and 7.4 (First Back Arc Between Consecutive Containers). ∎